OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x2 - 6x + 2 = 0 A. {0.4, 5.6} B. {0.5, 4.8} C. {1.4, 2.4} D. no real number solution

OpenStudy (anonymous):

That isn't a perfect square.

OpenStudy (hoblos):

A do you want me to post the explanation ?

OpenStudy (anonymous):

yes......

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{\begin{array}{l} {x^2} - 6x + 2 = 0\\ - b \pm \sqrt {{b^2} - 4ac} \end{array}}{{2a}}\\ \frac{{6 \pm \sqrt {{{( - 6)}^2} - 4(1 \cdot 2)} }}{{2 \cdot 1}}\\ = \frac{{6 \pm \sqrt {28} }}{2}\\ = {\rm{5}}{\rm{.65,0}}{\rm{.35}} \end{array}\]

OpenStudy (hoblos):

fewscrewsmissing did it ;)

OpenStudy (anonymous):

I don't understand how you get 2 values......

OpenStudy (anonymous):

That's the roots of the equation - plug either into the polynomial, and it'll equate to 0. (NB: I've rounded here, so THOSE particular numbers won't be 0; but the actual explicit solution).

OpenStudy (hoblos):

pratu043 notice the +/-

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

plus or minus

OpenStudy (anonymous):

pratu043 the "quadratic formula" is well worth remembering - precisely for when factoring is difficult.